# Rethinking Metrics and Benchmarks of Video Anomaly Detection

### v1

@@ 3.3 Current Benchmarks Struggle to Evaluate Scene Overfitting
Normal Abnormal AUC FAR0.5
train set 01-06 07-10 - -
easy test set 11,12,13 11,12,13 94.55 0.38
hard test set 07-10 01-06 16.35 44.10
Existing VAD models trained on web-sourced large-scale datasets [13, 15, 14] may risk overfitting to training scenes, yet current benchmarks cannot fully evaluate the overfitting, because they comprise scenes different from training data and fail to evaluate whether models retain performance on scenes that closely resemble training conditions. For instance, the production scenario may be a crowded grocery store with a checkout counter on the right, which is included in the training data, while existing test data only comprises videos of mobile phone stores and clothing retailers.
Therefore, we conduct a preliminary experiment by re-splitting the ShanghaiTech dataset [28] as shown in Table 3. ShanghaiTech is a manually recorded dataset with normal and abnormal videos across 13 scenes. We first construct a training set and a easy test set to simulate the case of web-sourced datasets like UCF-Crime, where each scene in the training set contains exclusively normal or exclusively abnormal videos, with test scenes being distinct from those in the training set. Subsequently, we construct a hard test set containing identical scenes to the training set but with reversed video normality in the same scenes. We train PEL [9] on the training set, evaluate on both test sets with the best checkpoint. As shown in Table 3, while the model demonstrates satisfactory performance on the easy test set, which simulates existing benchmarks, it significantly deteriorates on the hard test set, resulting in a sharp AUC drop to 16% with a high 44% FAR, indicating severe scene overfitting.
4 Proposed Metrics and Benchmarks
4.1 Annotation Averaged AUC and AP Metrics
To mitigate the impact of bias in single-round annotation, one straightforward solution is to standardize the annotation process, particularly regarding granularity and ending criteria. However, this is challenging given the complexity and diversity of anomalies. Alternative approaches include employing soft labels or aggregating via voting mechanisms, but they either fail to be compatible with the existing metrics or cause information loss. Consequently, we propose to calculate AUC/AP separately across annotations and use the averages as more reliable metrics, which avoids the aforementioned drawbacks. Formally,
$\text{AUC}_{\text{avg}}=\sum_{r=0}^{R}\text{AUC}(\hat{\mathcal{Y}},\mathcal{Y}_{r})/R,\quad\text{AP}_{\text{avg}}=\sum_{r=0}^{R}\text{AP}(\hat{\mathcal{Y}},\mathcal{Y}_{r})/R,$ (1)
where $\hat{\mathcal{Y}}$ is the model predictions, $\mathcal{Y}_{r}$ is the $r$-th round annotation, and $R$ denotes the number of annotation rounds.
4.2 Latency-aware AP Metric
Motivated by the latency sensitivity of real-world anomaly detection systems, we introduce Latency-aware Average Precision (LaAP) - a novel evaluation metric that expands the standard AP metric through a time-decaying weighting mechanism, explicitly rewarding earlier predictions. Formally, LaAP is defined as the area under the Precision-LaRecall curve under all possible thresholds $\tau$:
@@ 4.2 Latency-aware AP Metric
where $\hat{y}$ denotes the model prediction scores, the superscript $j$ denotes $j$-th video in the dataset, the subscript $i$ denotes $i$-th frame in the video, $M_{j}$ denotes number of frames in the $j$-th video, and $\overline{Y}^{j}$ is the binary labels.
@@
In practice, LaAP assumes that only one abnormal event exists in a video due to the sparsity of anomalies. In existing datasets, almost all data satisfy this assumption except for the XD-Violence dataset, which has many heavily edited videos and contains multiple logically unrelated events within a single video. Consequently, we do not compute the LaAP metric on the XD-Violence dataset. For multi-round annotations, we select the median of the first abnormal frames and the latest abnormal frames across different annotations as the starting and ending times. This also mitigates the single annotation bias issue discussed in Section 3.1. Empirically, we set the parameter of LaAP as follows: $\alpha=2,\beta=7,\phi=16$, where $\alpha$ halves the weight of the next point, $\beta$ builds a scoring function that is close to 1 or 0 at endpoints, and $\phi$ ensures that the score is only calculated once in a short period.
4.3 Video-Diffusion-Generated Hard Normal Benchmarks
@@
Specifically, we first manually select scene keyframes from abnormal videos in the UCF-Crime and MSAD training sets. We prioritize frames with minimal subject interference, absence of motion blur, and high visual quality, which is better for diffusion models to generate high-quality videos. Only one keyframe is selected per video. Due to computational efficiency constraints, we sample a subset of videos for generation. Then, we use image-to-video diffusion models with instance-specific manual prompts to generate normal videos preserving identical scenes. Finally, we check all generated videos and re-generate videos that exhibit low quality or artifacts. For some videos, multiple normal samples are synthesized, yielding 100 hard normal samples for UCF-HN and 67 hard normal samples for MSAD-HN, which last from 10 to 16 seconds. Due to the complexity of abnormal actions and inability of generating multi-shot videos, we only generate normal videos and exclude XD-Violence. We provide both quantitative and qualitative validation of the realism of generated videos in the following experiments.
To perform benchmarking, the model should be trained on UCF-Crime or MSAD, and then be tested on the corresponding hard normal set. Since all test videos are normal, the false alarm rate (FAR) reflects the degree of overfitting. More details and visualizations of the benchmarks are in the supplementary Section F.
5 Experiments
5.1 Baselines
We reproduced 10 SOTA VAD methods spanning four methodological groups: 1) Traditional methods (RTFM [3], MGFN [4], UR-DMU [5]); 2) Text-enhanced methods (PEL [9], VadCLIP [10], TEVAD [7], CLIP-TSA [8]); 3) Zero-shot methods (LAVAD [31]); 4) Extra supervision methods (GlanceVAD [32], HolmesVAU [33]). With priority given to officially released weights, we retrained unavailable models using their original implementations. Details of baselines are provided in the supplementary Section D.
Figure 6: Correlation analysis between LaAP and AP.
Figure 7: Visualization of different score modification methods.
Table 4: Comparisons between AUC, AP, and LaAP under different score modification methods. Our proposed LaAP is able to effectively reward earlier detections.
Model Modification AUC AP LaAP
VadCLIP Desc. 0.900 0.486 0.783
Ori. 0.900 0.486 0.740
Asc. 0.900 0.486 0.675
PEL Desc. 0.886 0.439 0.744
Ori. 0.886 0.439 0.727
Asc. 0.886 0.439 0.645
Figure 8: Visualizations of our proposed synthetic benchmarks.
5.2 Reliability of LaAP
In Figure 6, we plot the AP and LaAP values and calculate the Spearman’s Rank-Order Correlation Coefficient (SROCC) during the training of VadCLIP, which shows strong correlation between LaAP and AP. To investigate the advantages of LaAP, we conduct a comparative analysis by modifying predicted scores from VadCLIP [10] and PEL [9] on UCF-Crime [13]. As illustrated in Figure 7, we simulate early/late detection through temporal reordering of anomaly scores within abnormal segments while preserving original score magnitudes. The Desc. makes the detection earlier by sorting scores in descending order, while Asc. achieves the opposite effect through ascending order.
Table 5: Comparisons between UBNormal, UCF-HN and MSAD-HN with FID [49] and FVD [50] metrics.
Datasets FID FVD
UBNormal [27] 202.6 698.0
UCF-HN 149.5 530.2
MSAD-HN 140.4 461.0
Table 4 demonstrates LaAP’s latency awareness: While AUC/AP remain invariant, our LaAP effectively rewards earlier detections through its temporal weighting mechanism.
5.3 Realism of UCF-HN and MSAD-HN
To validate the realism of our synthetic data, Figure 8(a) visually compares an original anomaly video (showing “Abuse”) with its generated normal counterpart, where violent actions are replaced by natural walking while maintaining perfect background coherence. Figure 8(b) further analyzes the feature-space relationships: Our hard negative samples demonstrate closer proximity to abnormal videos than original normal videos. We further compute Fréchet Inception Distance (FID) [49] and Fréchet Video Distance (FVD) [50] between UBNormal [27] and our method on real-world surveillance videos, which comprise videos of original test sets in UCF-Crime and MSAD. As in Table 5, significantly lower values demonstrate the superior realism of our benchmarks.
5.4 Benchmarking Existing Methods
We report the performance of SOTA VAD methods with our proposed averaged AUC/AP, LaAP and hard normal benchmarks in Tables 6,7. From Table 6, we observe that using averaged instead of original AUC/AP values significantly impacts method rankings, demonstrating that metric reliability is affected by annotator disagreements. This effect is most pronounced on UCF-Crime, while being relatively minor on XD-Violence and MSAD, which aligns with our observations of annotation discrepancy discussed in Section 3.1. Furthermore, comparative analysis of LaAP metrics reveals that models with comparable AUC/AP may exhibit substantial differences in LaAP. For instance, on UCF-Crime, MGFN underperforms other methods in both AUC and AP metrics but achieves the highest LaAP among weakly-supervised approaches. Similarly on MSAD, UR-DMU and VadCLIP demonstrate identical AUC performance yet exhibit significant LaAP discrepancy. Experiments also confirm that methods incorporating additional supervisory signals achieve consistent improvements across all metrics.
From Table 7, while most methods maintain low FAR ($\leq 10\%$) on the original test set, they suffer dramatic performance degradation on our proposed hard normal benchmarks. We observe a 42% average increase at 0.5 threshold, and 36% at 0.8 threshold, with some methods exceeding 70% FAR. This reveals severe scene overfitting in existing approaches. Comparative studies of supervision strategies using the same UR-DMU backbone reveal that clearer supervisory signals can mitigate overfitting risks. Although current anomaly videos in the training sets contain both normal and abnormal intervals, video-level weak supervision struggles to guide models in learning true anomaly patterns.
@@ 5.4 Benchmarking Existing Methods
Table 6: Comparisons between various state-of-the-art VAD models with original AUC/AP (Orig.), our proposed averaged AUC/AP (Avg.), and our proposed LaAP metric. Note that the metrics on UCF-Crime exclude three categories as discussed in Section 3.1. The scores are colored according to the rank. $\ast$ denotes zero-shot methods. $\dagger$ denotes methods with extra supervision signals.
@@
AUC AP AUC AP AUC AP
Model Orig. Avg. Orig. Avg. LaAP Orig. Avg. Orig. Avg. Orig. Avg. Orig. Avg. LaAP
@@
RTFM [3] 0.835 0.859 0.239 0.388 0.480 - - - - 0.868 0.867 0.662 0.645 0.713
MGFN [4] 0.820 0.863 0.216 0.384 0.761 - - - - 0.850 0.846 0.633 0.607 0.664
UR-DMU [5] 0.887 0.876 0.275 0.442 0.693 0.940 0.941 0.817 0.841 0.878 0.869 0.697 0.663 0.729
PEL [9] 0.890 0.886 0.305 0.439 0.727 0.949 0.945 0.855 0.861 0.873 0.869 0.676 0.659 0.729
VadCLIP [10] 0.889 0.900 0.254 0.486 0.740 0.947 0.952 0.845 0.870 0.878 0.865 0.597 0.541 0.665
TEVAD [7] 0.874 0.861 0.265 0.408 0.599 0.923 0.913 0.798 0.814 - - - - -
CLIP-TSA [8] 0.900 0.913 0.272 0.470 0.695 0.932 0.933 0.772 0.807 - - - - -
LAVAD [31] 0.856 0.842 0.336 0.446 0.592 0.854 0.859 0.620 0.687 - - - - -
GlanceVAD [32] 0.937 0.906 0.484 0.558 0.818 0.959 0.955 0.892 0.891 - - - - -
HolmesVAU [33] 0.893 0.912 0.321 0.547 0.800 - - - - - - - - -
Table 7: Performance comparison of SOTA VAD methods on UCF-Crime/MSAD and our proposed hard-normal variants (UCF-HN/MSAD-HN) using FARτ metric, where $\tau$ is the threshold. The scores are colored according to the value.
@@
UR-DMU [33] Full 0.005 0.000 0.136 $0$
6 Conclusion
In this paper, we reveal critical limitations in current metrics and benchmarks of VAD through systematic analysis. To address these limitations, we propose new metrics for reliable and latency-sensitive evaluation (i.e., averaged AUC/AP and LaAP). We also construct new benchmarks to evaluate the model generalizability (i.e., UCF-HN and MSAD-HN). We benchmark various SOTA VAD methods according to our new evaluation methods, providing novel perspectives on evaluating VAD algorithms.
Appendix A Limitation and Future Work
Although this study proposes two metrics for algorithm evaluation, their application remains confined to temporal localization and binary classification tasks. The current method does not extend to more granular evaluations such as spatial localization or multiclass classification, which could be developed in future research to establish a comprehensive evaluation system building upon the present work.
This research innovatively employs Artificial Intelligence-Generated Content (AIGC) technology to propose a method for evaluating model generalizability. However, our analysis reveals limitations in existing video generation methods, particularly in handling complex motion, small objects, temporal coherence, and prompt automation. Notwithstanding these challenges, we think that synthetically generated data exhibits substantial advantages over alternative data sources, as further elaborated in Section F.4. Subsequent studies could focus on developing scalable pipelines for producing high-quality synthetic video anomaly datasets.
Appendix B Open-source
The annotations files, codes to reproduce main results, notebooks, codes for LaAP calculation, and raw videos of proposed benchmarks will be provided.
Appendix C Details of Re-annotated Datasets
C.1 Selected Datasets
UCF-Crime [13] is the first large-scale and most commonly used VAD dataset of weak supervision, which comprises 1,900 surveillance videos from the web, covering 13 crime-related categories. This dataset employed 10 human annotators and took several months, with each video being assigned to multiple annotators and the final results averaged. However, the authors did not explicitly elaborate on the specific annotation method or the approach of the averaging process.
XD-Violence [15] is another large-scale dataset with 4,754 videos with an audio channel, covering 6 violence-related categories. The videos in XD-Violence may be labeled with more than one anomaly class. Its videos mainly come from movies, news, and user-uploaded videos on the web, which inevitably involve a lot of editing. XD-Violence employs the same annotation protocol as UCF-Crime but still provides only a single round of annotation without specifying the averaging process.
MSAD [14] is a recently proposed multi-scenario dataset, which is collected across 14 distinct scenarios, covering 11 base categories. The anomaly categories comprise both human-related and non-human-related anomalies, including some less serious anomalies (e.g., people falling). Moreover, MSAD provides videos of higher resolution. The authors of MSAD have not mentioned the details of the annotation process.
C.2 Details of Re-annotating
We re-annotated 2 rounds for UCF-Crime, 3 rounds for XD-Violence, and 3 rounds for MSAD. Each annotation round was implemented by a single annotator, with distinct annotators assigned to different rounds of annotation for the same dataset. During the annotation process, temporal segment annotations were applied to abnormal videos in the test set. Annotators were provided with the anomaly categories and definitions provided by each dataset, alongside explicit instructions that each video contained at least one abnormal segment. We used the Vidat annotation tool  [^2] https://github.com/anucvml/vidat. Each video took an average of 1 minute to annotate.
Especially, we conducted multi-class re-annotation for all 950 abnormal videos in UCF-Crime, with annotators instructed to assign each video to one of 14 predefined categories (including the normal class). Notably, although only abnormal videos were subjected to this relabeling, annotators were not informed of this constraint. Furthermore, these annotators constitute a distinct group from those involved in the temporal segment annotation, thereby avoiding potential bias.
C.3 Visualizations of Annotation Discrepancies
Figure A shows several videos that our annotators considered normal. In the first example, an individual placing clothes into a bag and exiting the frame could represent shoplifting behavior or a legitimate transaction occurring outside the camera’s view. The second case depicts a group of individuals carrying objects from a residence, with no visual evidence of illegal entry. The third instance involves an individual taking boxes from a doorway, where the viewer doesn’t know if he is the owner. Although these videos do have corresponding criminal behaviors, it is impossible to judge according to the video without prior knowledge.
We also provide visualizations of multiple forms of annotation discrepancies across three datasets in Figure B, validating the motivation of using averaged AUC/AP.
Figure A: Examples of multiclass annotation discrepancy in UCF-Crime. The abnormal events in these examples are hard to judge without further information.
(a) The start time of the anomaly is more consistent than the end time.
(b) Different annotation granularity in a surveillance video.
(c) Different annotation granularity in a heavily edited video.
(d) Multiple independent abnormal events, where the start time of the anomaly is more consensual.
(e) Example of low discrepancy where the entire video is abnormal.
(f) Example of low discrepancy where the boundary of the abnormal event is clear.
Figure B: Examples of multiple forms of annotation discrepancies.
Appendix D Details of Baselines

### v2

@@ III-C Current Benchmarks Struggle to Evaluate Scene Overfitting
Normal Abnormal PEL [9] VadCLIP [10]
@@
train set 01–06 07–10
easy test set 11,12,13 11,12,13 94.55 0.38 77.13 17.53
hard test set 07–10 01–06 16.35 44.10 36.46 53.86
Therefore, we conduct a preliminary experiment by re-splitting the ShanghaiTech dataset [31] as shown in Table III. ShanghaiTech is a manually recorded dataset with normal and abnormal videos across 13 scenes. We first construct a training set and a easy test set to simulate the case of web-sourced datasets, where each scene in the training set contains exclusively normal or exclusively abnormal videos, with test scenes being distinct from those in the training set. Subsequently, we construct a hard test set containing identical scenes to the training set but with reversed video normality in the same scenes. We train two SOTA models [9, 10] on the training set, evaluate on both test sets with the best checkpoint. As shown in Table III, while the models demonstrate satisfactory performance on the easy test set, which simulates existing benchmarks, they significantly deteriorate on the hard test set, resulting in a sharp AUC dropping (16%/36%) with a high FAR (44%/54%), indicating severe scene overfitting.
Finding 4. Although existing weakly/fully supervised methods achieve strong results on easy test sets with disjoint scenes, they have the potential risk of scene overfitting.
IV Proposed Metrics and Benchmarks
IV-A Preliminaries
We introduce the calculations of the traditional VAD metrics AUC and AP as follows. First, the predicted probabilities $\hat{y}_{i}^{j}$ and binary ground truths $y_{i}^{j}$ of all videos are concatenated into two long sequences:
$\mathbf{\hat{y}}$ $=[\hat{y}_{1}^{1},\hat{y}_{2}^{1},\dots,\hat{y}_{M_{1}}^{1},\hat{y}_{1}^{2},\dots,\hat{y}_{M_{V}}^{V}]\in{[0,1]}^{N},$ (1)
$\mathbf{y}$ $=[y_{1}^{1},y_{2}^{1},\dots,y_{M_{1}}^{1},y_{1}^{2},\dots,y_{M_{V}}^{V}]\in{\{0,1\}}^{N},$ (2)
where $j$ indexes videos and $i$ indexes frames within each video. $M_{j}$ is the frame count of video $j$, and $V$ is the total number of videos. Then, given a threshold $\tau$, the items of the confusion matrix are defined as:
$\mathrm{TP}$ $=\textstyle\sum_{j=1}^{V}\sum_{i=1}^{M_{j}}\mathds{1}(\hat{y}_{i}^{j}\geq\tau)\,y_{i}^{j},$ (3)
$\mathrm{FP}$ $=\textstyle\sum_{j=1}^{V}\sum_{i=1}^{M_{j}}\mathds{1}(\hat{y}_{i}^{j}\geq\tau)\,(1-y_{i}^{j}),$ (4)
$\mathrm{FN}$ $=\textstyle\sum_{j=1}^{V}\sum_{i=1}^{M_{j}}\mathds{1}(\hat{y}_{i}^{j}<\tau)\,y_{i}^{j},$ (5)
$\mathrm{TN}$ $=\textstyle\sum_{j=1}^{V}\sum_{i=1}^{M_{j}}\mathds{1}(\hat{y}_{i}^{j}<\tau)\,(1-y_{i}^{j}).$ (6)
We define a general area-under-curve operator $\Psi(X,Y)$ as:
$\Psi(X,Y)=\textstyle\int^{1}_{0}Y(\tau)\ dX(\tau),$ (7)
where $X,Y$ are two functions of $\tau$, called the integrator and the integrand, respectively. AUC/AP are defined as:
$\mathrm{R}=\mathrm{TPR}$ $=\mathrm{TP}/(\mathrm{TP}+\mathrm{FN}),$ (8)
$\mathrm{P}$ $=\mathrm{TP}/(\mathrm{TP}+\mathrm{FP}),$ (9)
$\mathrm{FPR}$ $=\mathrm{FP}/(\mathrm{FP}+\mathrm{TN}),$ (10)
$\mathrm{AUC}=\Psi(\mathrm{FPR},$ $\mathrm{TPR}),\quad\mathrm{AP}=\Psi(\mathrm{R},\mathrm{P}).$ (11)
IV-B Probabilistic AUC and AP Metrics
To mitigate the impact of bias, we ask: can standardizing the annotation guideline solve the problem? We argue it cannot, primarily for two reasons: 1) the intrinsic complexity and diversity of anomalies make it costly to train annotators, especially given a large number of categories; 2) more critically, any fixed guideline induces a specific dataset bias. Consequently, an algorithm’s poor performance may stem from a misalignment with this bias rather than a genuine lack of capability. Therefore, the concept of a deterministic ground truth is inadequate. Instead, a more robust approach is to define the ground truth as a distribution and approximate it by utilizing multiple annotations.
We explored three potential approaches of utilizing multiple annotations. One straightforward approach is to aggregate annotations via a voting mechanism. Nevertheless, using hard labels incurs a loss of information and cannot avoid bias. We also tried calculate AUC/AP separately across annotations and use the average as a more reliable metric. However, a significant limitation is the lack of a clear upper bound, as a prediction would need to perfectly match all annotations to achieve a score of 1, which is impossible because the annotations are different from each other. Consequently, inspired by [47], we adopted the approach of first averaging multi-round annotations to create a soft label, and then extend the traditional AUC/AP to probabilistic AUC/AP (i.e., ProbAUC, ProbAP) to utilize the soft labels, which avoids the aforementioned drawbacks.
Figure 5: Comparison between vanilla ROC/P-R curve and probabilistic ROC/P-R curve.
Formally, ProbAUC and ProbAP extend the conventional evaluation framework by substituting the binary ground truth vector $\mathbf{y}$ with a probabilistic ground truth vector $\tilde{\mathbf{y}}$:
$\tilde{\mathbf{y}}=[\tilde{y}_{1}^{1},\tilde{y}_{2}^{1},\dots,\tilde{y}_{M_{1}}^{1},\tilde{y}_{1}^{2},\dots,\tilde{y}_{M_{V}}^{V}]\in{\{0,1\}}^{N},$ (12)
Subsequently the components of the confusion matrix and the related base metrics can be derived in a manner analogous to Eqs. (3)-(6),(8)-(10). However, the introduction of probabilistic labels alters the effective range of the performance metric. As illustrated in Fig. 5, the area under the curve corresponding to the best classifier is no longer one, while that of the worst classifier deviates from zero. Accordingly, we introduce ProbAUC and ProbAP with a normalized definition:
$\mathrm{ProbAUC}$ $=\frac{\Psi(\mathrm{FPR},\mathrm{TPR})-\Psi(\mathrm{FPR^{{\color[rgb]{0,0.5234375,0.5234375}w}}},\mathrm{TPR^{{\color[rgb]{0,0.5234375,0.5234375}w}}})}{\Psi(\mathrm{FPR^{{\color[rgb]{0.78125,0.3203125,0.1640625}b}}},\mathrm{TPR^{{\color[rgb]{0.78125,0.3203125,0.1640625}b}}})-\Psi(\mathrm{FPR^{{\color[rgb]{0,0.5234375,0.5234375}w}}},\mathrm{TPR^{{\color[rgb]{0,0.5234375,0.5234375}w}}})},$ (13)
$\mathrm{ProbAP}$ $=\frac{\Psi(\mathrm{R},\mathrm{P})-\Psi(\mathrm{R^{{\color[rgb]{0,0.5234375,0.5234375}w}}},\mathrm{P^{{\color[rgb]{0,0.5234375,0.5234375}w}}})}{\Psi(\mathrm{R^{{\color[rgb]{0.78125,0.3203125,0.1640625}b}}},\mathrm{P^{{\color[rgb]{0.78125,0.3203125,0.1640625}b}}})-\Psi(\mathrm{R^{{\color[rgb]{0,0.5234375,0.5234375}w}}},\mathrm{P^{{\color[rgb]{0,0.5234375,0.5234375}w}}})},$ (14)
where the superscript $w$ denotes the worst classifier and $b$ denotes the best classifier. From a geometric perspective, the metric can be interpreted as:
[…]
Notably, when $\tilde{\mathbf{y}}$ is reduced to a hard label, ProbAUC and ProbAP degenerate to the conventional AUC and AP, i.e., $\Psi(\mathrm{FPR^{{\color[rgb]{0,0.5234375,0.5234375}w}}},\mathrm{TPR^{{\color[rgb]{0,0.5234375,0.5234375}w}}})=\Psi(\mathrm{R^{{\color[rgb]{0,0.5234375,0.5234375}w}}},\mathrm{P^{{\color[rgb]{0,0.5234375,0.5234375}w}}})=0$ and $\Psi(\mathrm{FPR^{{\color[rgb]{0.78125,0.3203125,0.1640625}b}}},\mathrm{TPR^{{\color[rgb]{0.78125,0.3203125,0.1640625}b}}})=\Psi(\mathrm{R^{{\color[rgb]{0.78125,0.3203125,0.1640625}b}}},\mathrm{P^{{\color[rgb]{0.78125,0.3203125,0.1640625}b}}})=1$.
IV-C Latency-aware AP Metric
Motivated by the latency sensitivity of real-world anomaly detection systems, we introduce Latency-aware Average Precision (LaAP), which expands the standard AP metric through a time-decaying weighting mechanism, explicitly rewarding earlier predictions. Formally, LaAP is defined as the area under the Precision-LaRecall curve under all possible thresholds $\tau$:
$\mathrm{LaAP}=\int_{0}^{1}\text{LaRecall}(\tau)\ d\mathrm{P}(\tau),$ (16)
@@ IV-C Latency-aware AP Metric
where $\overline{Y}^{j}$ is the binary predictions of the j-th video.
@@
Figure 6: Visualization of the computational steps of LaRecall.
In practice, LaAP assumes that only one abnormal event exists in a video due to the sparsity of anomalies. In existing datasets, almost all data satisfy this assumption except for the XD-Violence dataset, which has many heavily edited videos and contains multiple logically unrelated events within a single video. Consequently, we do not compute the LaAP metric on the XD-Violence dataset. For future datasets containing multiple independent anomalies in a single video, LaAP can be applied by splitting the video into segments. For multi-round annotations, we select the median of the first abnormal frames and the latest abnormal frames across different annotations as the starting and ending times. This also mitigates the single annotation bias issue discussed in Section III-A.
We provide guidelines and default settings for three parameters as follows:
  • $\phi$: This parameter is to avoid redundant scoring, which is a positive integer representing the sampling interval. It should remain fixed at 16 for most cases.
  • $\alpha$: This parameter controls the relative importance of sampling points, which should be a positive number greater than 1. A higher $\alpha$ reduces attention to later sampling points. With our default $\alpha=2$, the weights for points beyond six points decay below 1%. To prioritize earlier detections, $\alpha$ should be increased.
  • $\beta$: This parameter controls the steepness of the Sigmoid-like scoring curve, determining LaAP’s tolerance near event boundaries, which takes a positive number. As $\beta$ approaches 0, the curve exhibits more linearity. And larger values make it approximate a step function. With our default $\beta=7$, sampling points within the 34% interval preceding anomaly onset receive scores above 0.9. To allow for greater detection latency, $\beta$ can be increased.
IV-D Video-Diffusion-Generated Hard Normal Benchmarks
@@
TABLE IV: Comparisons between various state-of-the-art VAD models with original AUC/AP, our proposed ProbAUC/ProbAP, and our proposed LaAP metric. Note that the metrics on UCF-Crime exclude three categories as discussed in Section III-A. The scores are colored according to the rank. $\ast$ denotes zero-shot methods. $\dagger$ denotes methods with extra supervision signals.
@@
Model AUC ProbAUC AP ProbAP LaAP AUC ProbAUC AP ProbAP AUC ProbAUC AP ProbAP LaAP
@@
RTFM [3] 0.835 0.864 0.239 0.434 0.480 - - - - 0.868 0.870 0.662 0.665 0.713
MGFN [4] 0.820 0.868 0.216 0.428 0.761 - - - - 0.850 0.849 0.633 0.626 0.664
PEL [9] 0.890 0.875 0.305 0.495 0.727 0.949 0.944 0.855 0.856 0.873 0.873 0.676 0.683 0.729
UR-DMU [5] 0.887 0.889 0.275 0.489 0.693 0.940 0.950 0.817 0.881 0.878 0.873 0.697 0.678 0.729
VadCLIP [10] 0.889 0.908 0.254 0.542 0.740 0.947 0.956 0.845 0.885 0.878 0.869 0.597 0.557 0.665
TEVAD [7] 0.874 0.861 0.265 0.455 0.599 0.923 0.917 0.798 0.829 - - - - -
CLIP-TSA [8] 0.900 0.920 0.272 0.526 0.695 0.932 0.936 0.772 0.820 - - - - -
LAVAD [34] 0.856 0.847 0.336 0.491 0.592 0.854 0.865 0.620 0.700 - - - - -
GlanceVAD [35] 0.937 0.902 0.484 0.612 0.818 0.959 0.960 0.892 0.911 - - - - -
HolmesVAU [36] 0.893 0.921 0.321 0.610 0.800 - - - - - - - - -
(1) Select Keyframes: To obtain high-quality images, we manually extracted key frames from referenced abnormal videos in the training sets of UCF-Crime and MSAD. Due to the complexity of abnormal actions, we only generate normal videos. The extracted keyframes meet the following criteria: 1) No inclusion of abnormal events; 2) Absence of motion blur; 3) Minimal visible watermarks; 4) One keyframe per video.
(2) Generate Videos: We used image-to-video diffusion models (Vidu 2.0[32] and wan2.1_i2v_480p_14B[33]) with instance-specific manual prompts to generate normal videos preserving identical scenes. Each video was created by first generating a clip from the selected key frame, then producing an extended clip using the last frame of the first clip, followed by temporal concatenation. The generated videos last from 10 to 16 seconds. Manual prompt engineering ensured diverse event representations in generated videos.
(3) Recheck: During this phase, we re-generated videos exhibiting suboptimal quality while discarding frames with persistent synthesis artifacts. All retained videos were subjected to manual quality inspection to verify visual fidelity. This process yielded 100 hard normal samples for UCF-HN and 67 hard normal samples for MSAD-HN. These two benchmarks are intended to complement the test sets of UCF-Crime and MSAD rather than replacing them entirely. Compared to the numbers of normal test videos of these datasets (150 in UCF-Crime, 120 in MSAD), we consider our benchmarks’ scale sufficient for their purpose.
We provide both quantitative and qualitative validation of the realism of generated videos in the following experiments. To perform benchmarking, the model should be trained on UCF-Crime or MSAD, and then be tested on the corresponding hard normal set. Since all test videos are normal, the false alarm rate (FAR) reflects the degree of overfitting.
TABLE V: Performance comparison of SOTA VAD methods on UCF-Crime/MSAD and our proposed hard-normal variants (UCF-HN/MSAD-HN) using FARτ metric, where $\tau$ is the threshold. The scores are colored according to the value.
@@
UR-DMU [36] Full 0.005 0.000 0.136 0.000 - - - -
V Experiments
V-A Baselines